A Bibliography of Publications of Michel Wermelinger

Michel Wermelinger Universidade Nova de Lisboa Departamento de Informatica

2825 Monte da Caparica Portugal

Tel: ?n/a? FAX: ?n/a?

E-mail: mw@di.fct.unl.pt (Internet)

25 November 2011 Version 1.01

Abstract

This bibliography records publications of Michel Wermelinger.

Title word cross-reference

Abstract [Wer98b]. Acquisition [WL93]. Aéreas [Wer96]. AIPO [Wer90]. Algebra [WF98b]. Ambiente [Wer90]. Analysis [Wer98a]. ao [WLC92]. Apoio [WLC92]. Aquisição [WL91b]. Architectural [WF98b]. Architecture [Wer97b].

Base [WL91a]. baseada [WL91b]. based [WL93]. Basic [WL94]. Básica [Wer95c].

Canonicity [Wer97a]. Case [WF98b]. Chemical [Wer98b]. CIM [WB93]. Conceptuais [WL91b].

Conceptual [Wer91b, WL93]. Conhecimento [WL91b]. Connectors [WF97].

Dados [WL91a]. Different [Wer97a]. Dynamic [Wer97b].

Editor [Wer91c]. Efficiently [Wer93]. Estruturas [Wer95c]. Estudo [WLC92]. Evolution [Wer98b]. EXPRESS [Wer92].

Ferramenta [WL91b]. First [Wer95a]. First-Order [Wer95a]. Folhas [WLC92]. Formal [Wer98a]. Fundamentos [WLC92].

GET [Wer91c]. Grafos [WL91b]. Graph [Wer91c]. Graphical [EWC^{+}92]. Graphs [Wer91b]. GRIP [EWC^{+}92].

Hierarchic [Wer97b].

Implementation [Wer91b]. Incomplete [Wer91c].

Interface [CPPW92]. Introdução [Wer90].

Knowledge [WL93].

Linear [CPPW92, CPPW92]. Logic [Wer95a].

Machine [Wer98b]. Maximização [Wer96]. Mobile [WF97]. Mobility [WF98b]. Model [Wer97b]. Modeling [WB93]. Modelização [WL91a].

Notation [CPPW92]. Notes [Wer91a].

Order [Wer95a].

para [WL91b]. PEIRCE [Wer95b]. Perspective [Wer97a]. Posets [Wer93]. Prático [WLC92]. Programação [Wer90, WLC92]. Programs [WF97]. Proposal [Wer95b].

Reconfiguration [Wer97b]. Reconstruction [Wer92]. Reference [Wer91c]. Representação [WL91b]. Representation [WL93].

Searching [Wer93]. Some [Wer91a]. Specification [Wer98a]. Structures [Wer92, WL93]. Study [WF98b]. Synchronization [WF98b].

Teoria [Wer95c]. Theory [WL94]. Toolkit [WL93]. Tools [Wer91c]. Transient [WF98c].

using [Wer92].

Viagens [Wer96].

Windows [WL93].

X [WL93]. X-Windows [WL93].

References

Choi:1992:LLN

[CPPW92] Young Bae Choi, Maurice Pagnucco, Heather Pfeiffer, and Michel Wermelinger. LINEAR: Linear notation interface. In Gerard Ellis and Robert A. Levinson, editors, Proceedings of the First International Workshop on PEIRCE: A Conceptual Graphs Workbench, pages 25–28. Key Centre for Software Technology, Dept. of Computer Science, University of Queensland, St. Lucia, Queensland., Australia, July

                                                  10, 1992. URL http://www-ctp.
                                                  di.fct.unl.pt/~mw/pubs/1992/
                                                  peirce92-l.ps.gz. Available as

University of Queensland Technical Report 241.

Esch:1993:LLN

[EPW93] John Esch, Maurice Pagnucco, and Michel Wermelinger. LINEAR— linear notation interface. In Robert Levinson and Gerard Ellis, editors, Proceedings of the Second International Workshop on PEIRCE: A Conceptual Graphs Workbench, pages 16–24. Laval University, Québec City, Canada, August 7, 1993. URL http:

                                                  //www-ctp.di.fct.unl.pt/~mw/
                                                  pubs/1993/peirce93.ps.gz.

Eklund:1992:GGI

[EWC^{+}92] Peter Eklund, Michel Wermelinger, Keith Campbell, Harry Delugach, Robert Muehlbacher, and Heather Pfeiffer. GRIP: Graphical interface. In Gerard Ellis and Robert A. Levinson, editors, Proceedings of the First International Workshop on PEIRCE: A Conceptual Graphs Workbench, pages 29–31. Key Centre for Software Technology, Dept. of Computer Science, University of Queensland, St. Lucia, Queensland., Australia, July

10, 1992. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1992/
peirce92-g.ps.gz. Available as

University of Queensland Technical Report 241.

Wermelinger:1993:CSM

[WB93] Michel Wermelinger and Alex Bejan. Conceptual structures for modeling in CIM. In Conceptual Graphs for Knowledge Representation, volume 699 of Lecture Notes in Artificial Intelligence, pages 345–360. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1993. URL http:

//www-ctp.di.fct.unl.pt/~mw/
pubs/1993/iccs93.ps.gz. Pro-

ceedings of the First International Conference on Conceptual Structures.

Wermelinger:1990:AAI

[Wer90] Michel Wermelinger. AIPO:

Ambiente de Introdução à Programação. Technical Report RT-36/90-DI, Departamento de Informática, Universidade Nova de Lisboa, May 1990. Manual de referência da versão 1.41.

Wermelinger:1991:GSN

[Wer91a] Michel Wermelinger. GET— some notes on the implementation. Technical Report RT-4/91-CIUNL, Centro de Informática da Universidade Nova de Lisboa, January 1991. URL http:

                                        //www-ctp.di.fct.unl.pt/~mw/
                                        pubs/1991/get-impl.ps.gz.

Wermelinger:1991:GIC

[Wer91b] Michel Wermelinger. GET: An implementation of conceptual graphs. Technical Report RT-5/91-CIUNL, Centro de Informática da Universidade Nova de Lisboa, January 1991.

                                        URL http://www-ctp.di.fct.
                                        unl.pt/~mw/pubs/1991/get.ps.
                                        gz.

Wermelinger:1991:GGE

[Wer91c] Michel Wermelinger. GET: Graph Editor and Tools—the incomplete reference. Technical Report RT-3/91-CIUNL, Centro de Informática da Universidade Nova de Lisboa, January

                                        1991. URL http://www-ctp.
                                        di.fct.unl.pt/~mw/pubs/1991/
                                        get-ref.ps.gz.

Wermelinger:1992:REU

[Wer92] Michel Wermelinger. A reconstruction of express using conceptual structures. Term paper for the “Information Systems for CIM” course, December 1992.

                                        URL http://www-ctp.di.fct.
                                        unl.pt/~mw/pubs/1992/cim.ps.

gz. MSc in Computer Engineering.

Wermelinger:1993:SEP

[Wer93] Michel Wermelinger. Searching efficiently in posets. Term paper for the “Topics in Programming Techonology” course, March 1993.

URL http://www-ctp.di.fct.
unl.pt/~mw/pubs/1993/ttp.ps.

gz. MSc in Computer Engineering.

Wermelinger:1995:CGF

[Wer95a] Michel Wermelinger. Conceptual graphs and first-order logic. In Conceptual Structures: Applications, Implementation and Theory, volume 954 of Lecture Notes in Artificial Intelligence, pages 323–337. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo,

1995. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1995/

iccs95.ps.gz. Proceedings of the Third International Conference on Conceptual Structures.

Wermelinger:1995:CSL

[Wer95b] Michel Wermelinger. Conceptual structures linear notation: A proposal for PEIRCE. In Peter Eklund, editor, Proceedings of the Fourth International Workshop on PEIRCE, pages 13–24. ????, Santa Cruz, California, USA, August

18, 1995. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1995/
peirce95.ps.gz.

Wermelinger:1995:TBE

[Wer95c] Michel Wermelinger. Teoria Básica das estruturas conceptuais. Master’s thesis, Universidade Nova de Lisboa, January

                                        1995. URL http://www-ctp.
                                        di.fct.unl.pt/~mw/pubs/1995/
                                        tese.ps.gz.

Wermelinger:1996:MAV

[Wer96] Michel Wermelinger. Maximização de viagens Aéreas. Nota Técnica UNL-DI-4/96, Departamento de Informática, Universidade Nova de Lisboa, July

                                        1996. URL http://www-ctp.
                                        di.fct.unl.pt/~mw/pubs/1996/
                                        viagens.ps.gz.

Wermelinger:1997:DPC

[Wer97a] Michel Wermelinger. A different perspective on canonicity. In Proceedings of the Fifth International Conference on Conceptual Structures, volume 1257 of Lecture Notes in Artificial Intelligence, pages 110–124. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1997. URL http:

                                        //www-ctp.di.fct.unl.pt/~mw/
                                        pubs/1997/iccs97.ps.gz.

Wermelinger:1997:HAM

[Wer97b] Michel Wermelinger. A hierarchic architecture model for dynamic reconfiguration. In Proceedings of the Second International Workshop on Software Engineering for Parallel and Distributed Systems, pages 243–254. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997.

                                        URL http://www-ctp.di.fct.
                                        unl.pt/~mw/pubs/1997/pdse97.
                                        ps.gz.

Wermelinger:1998:FSA

[Wer98a] Michel Wermelinger. Formal specification and analysis of dynamic reconfiguration of software architecture. In Proceedings of ICSE’98, volume 2. ACM Press, New York, NY 10036, USA, April

1998. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1998/

icse98.ps.gz. To appear. To be presented at the Doctoral Symposium.

Wermelinger:1998:SAE

[Wer98b] Michel Wermelinger. Software architecture evolution and the chemical abstract machine. Submitted, January 1998. URL http:

//www-ctp.di.fct.unl.pt/~mw/
pubs/1998/evolution.ps.gz.

Wermelinger:1998:TCM

[Wer98c] Michel Wermelinger. Towards a chemical model for software architecture reconfiguration. In Proceedings of the 4th International Conference on Configurable Distributed Systems. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1998. To appear.

Wermelinger:1997:CMP

[WF97] Michel Wermelinger and José Luiz Fiadeiro. Connectors for mobile programs. Submitted to IEEE Transactions on Software Engineering. This paper is a revised version of [WF98a]., December

1997. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1997/
mobile.ps.gz.

Wermelinger:1998:CMP

[WF98a] Michel Wermelinger and José Fiadeiro. Connectors for mobile programs. Technical Report DI-FCUL TR-98-1, Department of Computer Science, University of Lisbon, January

1998. URL http://www-ctp.di.
fct.unl.pt/~mw/pubs/1998/di-
fcul-tr-98-1.ps.gz.

Wermelinger:1998:TAA

[WF98b] Michel Wermelinger and José Luiz Fiadeiro. Towards an algebra of architectural connectors: a case study on synchronization for mobility. In Proceedings of the 9th International Workshop on Software Specification and Design. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April

1998. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1998/
iwssd98.ps.gz. To appear.

Wermelinger:1998:TC

[WF98c] Michel Wermelinger and José Luiz Fiadeiro. Transient connectors. Submitted, February

1998. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1998/
mobility.ps.gz.

Wermelinger:1991:MAC

[WL91a] Michel Wermelinger and José Gabriel Lopes. Modelização conceptual de uma base de dados. Technical Note 10/91-CIUNL, Centro de Informática da Universidade Nova de Lisboa, December

1991. URL http://www-ctp.di.
fct.unl.pt/~mw/pubs/1991/bd.
ps.gz.

Wermelinger:1991:FPA

[WL91b] Michel Wermelinger and José Gabriel Lopes. Uma ferramenta para aquisição e representação de conhecimento baseada em grafos conceptuais. In Anais do VIII Simpósio Brasileiro de Inteligência Artificial, pages 287– 294. Sociedade Brasileira de Computação, 1991. URL http://www-

ctp.di.fct.unl.pt/~mw/pubs/
1991/sbia91.ps.gz.

Wermelinger:1993:XWT

[WL93] Michel Wermelinger and José Gabriel Lopes. An X-Windows toolkit for knowledge acquisition and representation based on Conceptual Structures. In Conceptual Structures: Theory and Implementation, volume 754 of Lecture Notes in Artificial Intelligence, pages 262–271. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK /

etc., 1993. URL http://www-ctp.
di.fct.unl.pt/~mw/pubs/1993/

awocg93.ps.gz. Proceedings of the Seventh Annual Workshop on Conceptual Graphs.

Wermelinger:1994:BCS

[WL94] Michel Wermelinger and José Gabriel Lopes. Basic conceptual structures theory. In Conceptual Structures: Current Practices — Proceedings of the Second International Conference on Conceptual Structures, volume 835 of Lecture Notes in Artificial Intelligence, pages 144– 159. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1994.

                                        URL http://www-ctp.di.fct.
                                        unl.pt/~mw/pubs/1994/iccs94.
                                        ps.gz.

Wermelinger:1992:FAA

[WLC92] Michel Wermelinger, Luís Seabra Lopes, and João C. Carrapiço. Folhas de apoio ao estudo prático de Fundamentos de Programação. Technical Note NT-3/92-DI, Departamento de Informática, Universidade Nova de Lisboa, June

                                        1992. URL http://www-ctp.di.
                                        fct.unl.pt/~mw/pubs/1992/fp.
                                        ps.gz.